OpenStudy (anonymous):

Write in standard form an equation of the line passing through the given point with the given slope. slope = 5/4 ; (0, 0)

OpenStudy (anonymous):

a. 4x + 5y = 0 c. 4x - 5y = 0 b. 5x + 4y = 0 d. 5x - 4y = 0